# Model Merging in LLMs, MLLMs, and Beyond: Methods, Theories, Applications and Opportunities

### v4

@@ 2.4 Theories and Analysis of Model Merging
In addition to designing various advanced methods in §2.2 and §2.3, the theoretical and effectiveness analysis of model merging is also crucial. Currently, there is limited work on the theoretical analysis of model merging. Based on the source of the models to be merged, the existing theoretical analysis can be roughly divided into three categories: (i) model merging of different checkpoints in the same training trajectory, (ii) model merging of different models fine-tuned on the same dataset, and (iii) model merging of different models fine-tuned on different datasets or tasks.
@@
First, some analyses target model merging on the single-trajectory training, usually referring to stochastic weighted average (SWA) or exponential moving average (EMA). For example, Jain et al. [75] theoretically proved that the excess risk of the EMA is an upper bound of a bias term and a variance term in the context of least squares regression. The bias term depends on the initialization state of the parameters and decreases exponentially with the number of iterations once the model starts averaging. The variance term depends on the noise covariance inherent in the data, which decays at a faster rate when model averaging is used [8]. Similarly, Rame et al. [149] applies bias-variance decomposition to the domain generalization setting to explain why model averaging improves out-of-distribution performance. In addition, Hardt et al. [56] provide a stability bound for SWA under convex assumptions, while Wang et al. [201] further establish generalization bounds analysis in both convex and nonconvex cases.
@@
Second, some studies explain the merging of multiple models with different hyperparameter fine-tuning for the same dataset in terms of connectivity and flatness of the loss landscape [98, 166, 202, 12]. Specifically, some works apply the theory of linear mode connectivity (LMC) [51, 41, 183] of neural networks to explain model merging. LMC reveals that neural network loss minima are not isolated points in the weight space. Recent studies [47, 131, 42, 242] have shown that two independent models, starting from the same pre-trained model and fine-tuned with different configurations, usually satisfy LMC. In other words, LMC is a general phenomenon that typically appears in fine-tuned models based on the ”pretraining-finetuning” paradigm, which is the current standard in the machine learning community. Therefore, performing weight alignment according to LMC provides a robust validity guarantee for model merging [5, 88]. On the other hand, other studies explain model merging from the perspective of a flatter loss landscape [100], arguing that merging multiple weights fine-tuned under different optimization configurations with the same data usually converges to a flat local minimum [45], thus revealing why model merging has better generalization [168, 73, 54, 232, 17].
Finally, an analysis by Ortiz-Jimenez et al. [136] is based on multiple models fine-tuned on different datasets, identifying weight disentanglement as a necessary precondition for effective model merging. More specifically, Ortiz-Jimenez et al. [136] provide theoretical and empirical analyses of the neural tangent kernel (NTK) and establish a compelling link between the task arithmetic [71] and the spectral properties of NTK.
3 Application of Model Merging in Foundation Models
@@
The emergence of foundation models, including large language models (LLMs), multimodal large language models (MLLMs), and image generative models, is a significant indicator of technological progress in the field of artificial intelligence in recent years. However, despite their advancements, these large models still face several challenges, such as generating harmful content in LLMs, MLLMs struggling with fusing information from different modalities, and the difficulty of producing mixed-style images in image generation models. Recent studies suggest that model merging techniques offer a promising solution to these inherent challenges in foundational models. Table 2 first briefly summarizes the application of model merging in foundational models. Then, §3.1, §3.2 and §3.3 provide a detailed discussion on how LLMs, MLLMs, and image generative models benefit from model merging, respectively.
@@
Table 2: A summary of the application of model merging techniques in foundation models.
@@
3.1 Model Merging in Large Language Models (LLMs)
@@
In recent years, large language models (LLMs), such as GPT-4 [2], Gemini [184], PaLM [25] and LLaMA [188], have made significant advancements and have been widely applied across various tasks. Despite their superhuman performance on most basic tasks, LLMs still face numerous challenges, including producing toxic content that violates laws or ethics, using unauthorized data during training, high training costs, and insufficient performance in specific domains. Model merging technology presents a promising opportunity to address these challenges.
3.1.1 Human Preference Alignment for LLMs
Humans often hold diverse opinions about aesthetics, politics, or fairness. When LLMs serve humans, different people have different expectations of the model, e.g., some expect LLMs to generate harmless responses, while others seek engaging and enjoyable interactions [151]. Consequently, the development of practical LLMs is generally divided into three stages, to generate responses that are more helpful, accurate, and safer [120]: Pre-training on a large amount of unsupervised data, supervised fine-tuning (SFT) on a small dataset with high-quality annotation, and interaction with humans to further optimize LLM alignment (e.g., direct preference optimization (DPO) [147] or reinforcement learning from human feedback (RLHF) [244]) with human preferences, rewards, or values.
Some works propose to achieve better, safer, or faster alignment of human preferences by model merging [238, 199, 13, 48, 141]. For example, ExPO [238] adds a task vector, constructed by a moderate model aligned using DPO or RLHF on a small amount of human preference data, to an unaligned SFT model. A more powerful aligned model can be directly obtained by setting a suitable merging coefficient. On the AlpacaEval 2.0 benchmark [110], fusing a model aligned on the 10%/20% preference data with an SFT model results in performance comparable to that of a model aligned on the full preference data. DogeRM [111] proposed merging the reward model with LLMs fine-tuned on different downstream domains to create domain-private reward models directly. Additionally, Lu et al. [120] propose an Online Merging Optimizer, that interpolates the gradient with the SFT model at each step of RLHF. This approach encourages RLHF to optimize toward reward maximization while preventing LLMs from forgetting general knowledge due to RLHF. Beyond preference alignment, several studies have examined the impact of model merging for secure alignment of LLMs [13, 225, 55, 57]. For example, Hammoud et al. [55] find that merging two security-aligned models could compromise security. Thus, they proposed explicitly including secure alignment as an optimization objective when constructing synthetic data for model merging.
In practice, users often have various combinations of preferences rather than a single preference. Training a model separately for each combination of preferences is unrealistic due to the infinite combinations and the high training costs. Therefore, some studies suggest combining models with different reward alignments to create a series of integrated aligned LLMs. For example, Rame et al. [151] and Jang et al. [78] propose Reward Soups and Personalized Soups, respectively, as efficient and flexible solutions for diverse rewards. Specifically, Rewarded Soups first trains an expert model for each reward and then linearly interpolates the weights of the experts to approximate the set of Pareto optimal solutions for various reward combinations. This approach is cost-effective, as it only requires training separate models for each reward to combine any variety of rewards.
3.1.2 Detoxifcation of LLMs
LLMs have been widely noted for issues related to untruthfulness and toxicity in various applications [66], such as insults, threats, and profanity in responses to certain questions. To address the potential security risks in the application of LLMs, flexible techniques are needed to reduce the generation of toxic text, essentially detoxifying LLMs. A straightforward solution is to collect additional non-toxic data to fine-tune LLMs [92]; however, this approach requires significant computing resources and may interfere with the general capabilities of LLMs. Alternatively, directly reducing the probability of potentially toxic words during the decoding stage requires additional guidance information [97]. Recent studies have shown that reducing the toxic data generation of LLMs through model merging is a simple and effective scheme [71, 234, 66, 37].
Task Arithmetic [71] negates the task vectors of GPT-2 model [146] fine-tuned on toxic data (Civil Comments [15]) and shows that this operation effectively reduces the proportion of data classified as ”toxic”, with little change in the fluency of the language on the control task (WikiText-103). Additionally, some parameter-efficient models steer the toxic behavior of LLMs by manipulating a small number of parameters. PEM [234] negates LoRA [65] (and (IA)3 [115]) modules trained on poisoning data to maintain language proficiency while reducing toxicity of language model output. Ethos [50] and Ext-Sub [66] point out that while the task vector on toxic data is factually wrong, it also contains correct information about language modeling and logical narrative skills. Therefore, Ext-Sub decomposes the toxic task vector into two orthogonal subspaces that represent general capability and destructive capability, respectively. Toxic knowledge is then eliminated by removing only the component representing the destructive ability from the LLM.
3.1.3 Knowledge Unlearning of LLMs
LLMs may inadvertently learn copyrighted material, raising significant legal and ethical concerns [1], and broader questions about responsible AI use [40]. In this context, the California Consumer Privacy Act [139] and the General Data Protection Regulations of the European Union [63] stipulate the right to data forgetting. The foundational model’s knowledge must be adapted to comply with these regulations. However, the cost of excluding copyrighted data for re-training from scratch is prohibitive. For instance, training a Llama-2-70B from scratch requires 1,720,320 GPU hours [189]. Traditional methods often use gradient ascent (GA) to achieve forgetting by fine-tuning the model using the GA algorithm on the specific data to be forgotten [187, 222]. Unfortunately, this approach typically catastrophically destroys other parts of the model’s knowledge. That is, forgetting specific knowledge also erases other knowledge that should be retained. Recently, many studies based on model merging techniques have demonstrated the potential to forget LLM-specific knowledge without harming other knowledge [71, 40, 66].
Unlike the GA-based approach, the model merging approach does not require additional data for other tasks to maintain old knowledge. To achieve forgetting, model merging typically incorporates a negatively fine-tuned model into the target model (i.e., the task-specific fine-tuned knowledge is subtracted from the target model). For example, Task Arithmetic [71] shows that negating task vectors degrade performance on specific tasks without substantial changes to the control tasks. Experiments demonstrate that model merging can forget the knowledge of the target task in a fine-tuned model without harming performance on control tasks. Similarly, Stable Sequential Unlearning (SSU) [40] extends this forgetting to the setting of sequential unlearning on LLMs, where different copyrighted content must be unlearned at different time steps. Knowledge forgetting can also forget samples that represent bad behavior during pretraining. For instance, FuseToForget [230] employs model merging as a debiasing tool to reduce privacy issues in language models. FLearning [135] first subtracts the parameters related to the data to be forgotten and then fine-tunes the parameters with new data to achieve accurate knowledge updates. SKU [119] explores the forgetting of harmful data in LLM, which is a two-stage scheme. Initially, harmful data (e.g., harmful question-answer pairs) is used to fine-tune the parameters corresponding to the location of harmful knowledge in the LLM (i.e., the task vector), and then the task vector is negated from the LLM to mitigate undesirable behavior in the LLM effectively. Generally, incorporating the opposite (anti-expert) task vectors into the pre-trained model can effectively accomplish the task of machine unlearning.
3.1.4 Faster Training of LLMs
Training LLMs requires numerous iterations on massive data, making the training process extremely expensive [89, 148]. For example, training LLAMA2-70B with 2T tokens required 1,720,320 GPU hours [113]. Methods to accelerate LLM training include mixed-precision training, continual retraining, and pipeline parallelism. An orthogonal approach is checkpoint merging in training trajectories, which offers a simple and effective means to either speed up LLM training or enhance training performance at the same cost.
The first type of works incorporate checkpoints in a single training trajectory during LLM training to accelerate model training. For instance, LAWA [89] demonstrated that merging checkpoints during intermediate stages of model training speeds up the process. For example, training a ResNet50 model on the ImageNet dataset reduced the training time by 68 GPU hours, and training a RoBERTa-Base model on the WikiText-103 dataset saved 30 GPU hours. Sanyal et al. [160] further showed that the combination of checkpoint averaging in pre-trained trajectories and a high learning rate contributes to faster convergence. Checkpoint Merging [113] comprehensively evaluates the effectiveness of model merging at different stages of the Baichuan2 [217] LLM model pre-training process. The second type of work involves combining existing models to create a more powerful initial model, thereby accelerating learning speed and improving accuracy on downstream tasks. For example, Fusing [23] and ColD Fusion [39] mixture of multiple existing fine-tuned models as base models and used for downstream task fine-tuning shows that this merged model outperforms the naive pre-trained model.
3.1.5 Combine the Capabilities of Expert LLMs
LLMs exhibit strong generalizability in general tasks, but often lack knowledge in specific vertical domains. Pretrained LLMs typically require fine-tuning within different corporations to become expert LLMs in various fields. Integrating the expertise of multiple specialists is particularly critical for solving more complex tasks. Research on model merging techniques indicates that a composite LLM can be created by combining the parameters of different expert LLMs [226, 6, 34, 240, 193, 194, 227, 171, 227, 99]. For example, Dekoninck et al. [34] demonstrate the ability to flexibly control text generation by merging multiple LLMs with different styles and applying personalized weighting. Robust Weight Signatures [16] proposes a robustness “patching” framework via model merging to enhance the overall robustness of the model against various naturally corrupted versions of clean data. In summary, model merging offers a straightforward and effective strategy for enhancing LLM’s capabilities.

### v5

@@ 3.1. Theoretical Analysis
In addition to designing various advanced methods in §2.2 and §2.3, the theoretical and effectiveness analysis of model merging is also crucial. Currently, there is limited work on the theoretical analysis of model merging. Based on the source of the models to be merged, the existing theoretical analysis can be roughly divided into three categories.
@@
(i) Single-Trajectory Model Averaging. Some analyses target model merging on the single-trajectory training, usually referring to stochastic weighted average (SWA) or exponential moving average (EMA). For example, Jain et al. (2018) theoretically proved that the excess risk of the EMA is an upper bound of a bias term and a variance term in the context of least squares regression. The bias term depends on the initialization state of the parameters and decreases exponentially with the number of iterations once the model starts averaging. The variance term depends on the noise covariance inherent in the data, which decays at a faster rate when model averaging is used (Arpit et al., 2022). Similarly, Rame et al. (2022) applies bias-variance decomposition to the domain generalization setting to explain why model averaging improves out-of-distribution performance. In addition, Hardt et al. (2016) provide a stability bound for SWA under convex assumptions, while Wang et al. (2024d) further establish generalization bounds analysis in both convex and nonconvex cases.
@@
(ii) Same-Task Merging. Some studies explain the merging of multiple models with different hyperparameter fine-tuning for the same dataset in terms of connectivity and flatness of the loss landscape (Kuditipudi et al., 2019; Simsek et al., 2021; Wang et al., 2022; Benton et al., 2021). Specifically, some works apply the theory of linear mode connectivity (LMC) (Garipov et al., 2018; Draxler et al., 2018; Tatro et al., 2020) of neural networks to explain model merging. Here, we first give a formal definition of LMC (Nagarajan and Kolter, 2019). Given a dataset $\small\mathcal{D}$ and two trained models $\small\Theta^{(i)}$ and $\Theta^{(j)}$ such that $\small\mathcal{L}(\Theta^{(i)})\approx\mathcal{L}(\Theta^{(j)})$ on $\mathcal{D}$, if there exists $\lambda\in[0,1]$ satisfying the following condition, we say that $\small\Theta^{(i)}$ and $\small\Theta^{(j)}$ are linearly connected in the loss landscape: $\small\mathcal{L}\left(\lambda\Theta^{(i)}+(1-\lambda)\Theta^{(j)}\right)\approx\mathcal{L}\left(\Theta^{(i)}\right)\approx\mathcal{L}\left(\Theta^{(j)}\right).$ LMC indicates that the loss minima of neural networks are not isolated points in weight space; from the perspective of LMC, these minima can be regarded as equivalent. Recent studies (Frankle et al., 2020; Nagarajan and Kolter, 2019; Entezari et al., 2022; Zhou et al., 2024a) have shown that two independent models, starting from the same pre-trained model and fine-tuned with different configurations, usually satisfy LMC. Therefore, performing weight alignment following LMC can transform multiple models to be merged into a parameter basin with higher merging feasibility, thereby providing a robust validity guarantee for model merging (Ainsworth et al., 2023; Jordan et al., 2023). On the other hand, other studies explain model merging from the perspective of a flatter loss landscape (Li et al., 2018), arguing that merging multiple weights fine-tuned under different optimization configurations with the same data usually converges to a flat local minimum (Foret et al., 2021), thus revealing why model merging has better generalization (Smith and Gashler, 2017; Izmailov et al., 2018; Gupta et al., 2020; Zhang et al., 2020; Cha et al., 2021).
(iii) Cross-Task Merging. An analysis by Ortiz-Jimenez et al. (2023) is based on multiple models fine-tuned on different datasets, identifying weight disentanglement as a necessary precondition for effective task vector based merging. More specifically, weight disentanglement is formally defined as follows. Consider a parametric function $\small f:\mathcal{X}\times\Theta\rightarrow\mathcal{Y}$. We say that $f$ is weight-disentangled if, for a set of task vectors $\small\{\tau_{t}\}_{t=1}^{T}$ and the corresponding datasets $\small\{\mathcal{D}_{t}\}_{t=1}^{T}$, the following holds: $\small f\left({x};\Theta^{(0)}+\sum_{t=1}^{T}\lambda_{t}{\tau}_{t}\right)=\sum_{t=1}^{T}g_{t}\left({x};\lambda_{t}{\tau}_{t}\right)+g_{0}({x}),\;t=1,2,\ldots,T,$ where $\small g_{t}({x};\lambda_{t}{\tau}_{t})=0$ if the sample $x$ does not belong to task $t$, i.e., $\small x\notin\mathcal{D}_{t}$, and $\small\small g_{0}({x})=0$ if $x$ does belong to task $t$. Intuitively, each task vector ${\tau}_{t}$ contributes only to its own task-specific component $g_{t}$, while $g_{0}$ captures the task-agnostic part. Ortiz-Jimenez et al. (2023) provide theoretical and empirical analyses of the neural tangent kernel (NTK) and establish a compelling link between the task arithmetic (Ilharco et al., 2023) and the spectral properties of NTK. From the NTK formulation, task vectors can be viewed as directions in the NTK feature space; when they are ”approximately orthogonal” under the corresponding metric, they do not hinder each other during merging. In addition, Zhou et al. (2025) establish a connection between task vectors and task gradients. Under standard full-batch gradient descent, the task vector obtained after one epoch of fine-tuning is exactly equal to the negative gradient of the loss scaled by the learning rate, i.e., $\small\tau_{t}=\eta\nabla\mathcal{L}_{t}(\Theta^{(0)})$. For multi-epoch fine-tuning, this relationship becomes approximate rather than exact. Wang et al. (2025c) use Approximate Kinematics Theory to derive an upper bound on the number of models that can be effectively merged. Once this threshold is exceeded, further merging no longer leads to performance improvements.
Summary and Discussion. Overall, the LMC perspective provides a direct explanation of why linear interpolation becomes reasonable once model weights are properly aligned. However, most existing analyses are conducted under the assumption of shared tasks or shared initialization, which limits their ability to account for the effectiveness of merging across heterogeneous tasks and modalities. In contrast, the connection between flat minima in the loss landscape and model generalization offers a plausible explanation for why weight averaging can improve out-of-distribution (OOD) performance, although a clear theoretical link between flatness and mergeability is still lacking. Finally, NTK-based theory explains how the similarity or approximate orthogonality between task vectors can mitigate conflicts during merging, but most of these results rely on wide-network limits and linearization assumptions, and their validity for large-scale LLM settings remains to be empirically verified. Taken together, these theoretical lines emphasize complementary yet distinct perspectives, and a unified theoretical framework for model merging has not yet emerged—an important direction for future research.
3.2. Empirical Comparisons
3.2.1. Benchmarks
To enable a fair comparison of different model merging methods, a number of benchmark-style works have been proposed. In Table 3, we present a comparison of existing model merging benchmarks. Specifically, MergeKit (Goddard et al., 2024a) is one of the earliest toolkits for merging model checkpoints. It combines multiple LLMs to improve performance and generality, and most of the merging strategies it includes are classic, easy-to-implement methods. FusionBench (Tang et al., 2025a) provides a model merging benchmark on visual classification tasks, but offers only limited support for LLM-based merging. Realistic Evaluation (Tam et al., 2024a) instead focuses on a benchmark for merging models on visual generation tasks. In recent years, more benchmarks have shifted their attention to LLMs. MergeBench (He et al., 2025) and Hitit et al. (2025) merge multiple domain-specialized LLMs, while Model-GLUE (Zhao et al., 2024b) defines a selection and aggregation protocol over a heterogeneous model zoo with different architectures and initializations. OptMerge (He et al., 2025) further targets the merging of multimodal large language models. In addition, several works analyze model merging from different perspectives. Merging at Scale (Yadav et al., 2024) and Merging Scaling Law (Wang et al., 2025b) investigate how the number of merged models affects performance, but mainly focus on classical merging methods. Mergenetic (Minut et al., 2025) instead studies the impact of evolutionary algorithms on merging performance.
Table 3. Comparison of existing model merging benchmarks. Columns indicate whether each benchmark includes vision models (image classification/generation), language models (standard NLP tasks), large language models (complex text generation), multimodal large language models (text + vision/audio), and whether the benchmark is open-sourced. Here, =yes and =no.
Evaluation Vision Model Text Model Large Language Model Multimodal Large Language Model Open-Source
MergeKit (Goddard et al., 2024a)
FusionBench (Tang et al., 2025a)
Realistic Evaluation (Tam et al., 2024a)
Merging at Scale (Yadav et al., 2024)
Model-GLUE (Zhao et al., 2024b)
MergeBench (He et al., 2025)
Mergenetic (Minut et al., 2025)
Merging Scaling Law (Wang et al., 2025b)
Systematic Study of Model Merging (Hitit et al., 2025)
OptMerge (Wei et al., 2025a)
(a) ViT-B/32
(b) ViT-L/14
(c) RoBERTa-Large
(d) Llama-3.1
(e) InternVL2.5
Figure 6. Accuracy of Various Model Merging Methods on Five Backbones: (a) ViT-B/32, (b) ViT-L/14, (c) RoBERTa-Large, (d) Llama-3.1-8B, (e) InternVL-2.5. Specifically, 8 visual classification tasks are merged on ViT-B/32 and ViT-L/14, 8 NLP tasks on RoBERTa-Large, 5 text generation tasks on Llama-3.1, and 5 multimodal tasks on InternVL-2.5.
3.2.2. Comparison of Results
To gain a more intuitive understanding of the performance differences between various model merging methods, we present the performance of common merging methods across different domains.
Visual Model Merging. Two architectures, ViT-B/32 and ViT-L/14, were selected, and their performance when merging 8 tasks was evaluated following the Fusionbench (Tang et al., 2025a) protocol. As shown in Figures 6(a) and 6(b), several key observations emerge: (1) Pre-trained models exhibit poor performance, as they lack task-specific capabilities tailored to downstream tasks. (2) Independent models (trained separately for each task) achieve the best performance due to the absence of cross-task interference. (3) Simple merging methods (e.g., Weight Averaging) result in significant performance degradation compared to independent models. (4) Weighted merging methods (e.g., RegMean, AdaMerging) and subspace merging methods (e.g., TSV-M, ISO-Merging) substantially improve the performance of merged models. (5) Optimization-based WUDI-Merging and post-hoc alignment method Surgery also demonstrates excellent performance. (6) The dynamic merging method WEMoE achieves performance closest to the upper bound. (7) Models with larger parameter scales are generally easier to merge (i.e., ViT-L/14 v.s. ViT-B/32).
Language Model Merging. Following Cheng et al. (2025), RoBERTa-Large models trained on eight GLUE tasks are merged, with the results reported in Figure 6(c). (1) The pretrained model and individually fine-tuned models still serve as the lower and upper performance bounds, respectively. (2) Subspace-based merging methods do not yield a significant performance change compared with Task Arithmetic. (3) In contrast, the optimization-based WUDI-Merging method achieves a clear performance improvement.
Large Language Model Merging. Following MergeBench (He et al., 2025), five domain-specialized LLMs are merged, with the results reported in Figure 6(d). (1) All merging methods lie between the pretrained model and the individually fine-tuned models, indicating that merging is beneficial but still incurs a certain degree of knowledge loss. (2) More sophisticated merging strategies, such as subspace-based or weighted merging, can even underperform the basic Task Arithmetic baseline. This may be due to the removal of neurons from the task vectors, which discards important information, or to the limited number of samples, which makes it difficult to reliably estimate parameter importance. In particular, DARE prunes neurons at random, leading to a high probability of removing important neurons and thus degrading performance.
Multimodal Large Language Model Merging. Following OptMerge (He et al., 2025), multimodal models trained on five tasks are merged, and the results are reported in Table 6(e). (1) Naive parameter averaging leads to severe performance degradation, even performing worse than the pretrained model. (2) Weighted merging achieves a substantial improvement over simple averaging, and sparse subspace methods such as DARE and TIES-Merging also perform well, whereas low-rank subspace approaches like TSV-M and ISO-Merging offer only moderate gains. (3) The optimization-based WUDI-Merging still achieves the best performance among all merging methods. (4) Nevertheless, all model merging approaches remain noticeably behind the individually fine-tuned models.
Summary and Discussion. These results suggest that simple model merging strategies typically deliver only moderate performance, whereas more advanced approaches—such as subspace-based or weighted merging—can further alleviate conflicts during merging and thus yield better accuracy. Optimization-based methods and dynamic merging schemes usually achieve the strongest results. However, all merging methods still fall short of the upper bound given by individual models.
4. Application of Model Merging in Foundation Models
@@
The emergence of foundation models, including large language models (LLMs), multimodal large language models (MLLMs), and image generative models, is a significant indicator of technological progress in the field of artificial intelligence in recent years. However, despite their advancements, these large models still face several challenges, such as generating harmful content in LLMs, MLLMs struggling with fusing information from different modalities, and the difficulty of producing mixed-style images in image generation models. Recent studies suggest that model merging techniques offer a promising solution to these inherent challenges in foundational models. Table 4 briefly summarizes the application of model merging in foundational models.
@@
Table 4. A summary of the application of model merging techniques in foundation models.
@@
4.1. Model Merging in Large Language Models (LLMs)
@@
In recent years, LLMs (such as GPT (Achiam et al., 2023), Gemini (Team et al., 2023), PaLM (Chowdhery et al., 2023) and LLaMA (Touvron et al., 2023a)) have made significant advancements and have been widely applied across various tasks. Despite their superhuman performance on most basic tasks, LLMs still face numerous challenges, including producing toxic content that violates laws or ethics, using unauthorized data during training, high training/inference/reasoning costs, and insufficient performance in specific domains. As shown in Figure 7, model merging technology presents a promising opportunity to address these challenges.